OpenStudy (spring98):

Two lines, A and B, are represented by the equations given below: Line A: x + y = 2 Line B: 2x + y = 4 Which statement is true about the solution to the set of equations? There are infinitely many solutions. There are two solutions. There is one solution. There is no solution.

OpenStudy (loser66):

1soln

OpenStudy (spring98):

How do you know?

OpenStudy (loser66):

good!!! they have different slope --> they intersect at some where. With 2 lines: only possibilities: 1) // (no solution) 2) exactly the same (many solutions) \3) 1 solution No way to get 2 solutions.

OpenStudy (loser66):

Algebraic way: \(\begin{cases}x+y=2\\2x+y=4\end{cases}\) subtract them together x = 2, hence y = 0

OpenStudy (spring98):

K that's better! That's why I need to know because some people give wrong answers.

OpenStudy (spring98):

Or help.

OpenStudy (loser66):

ok!! whatever, as long as you want to study, I am ok. Don't take the free answer. It's not good for the long run

OpenStudy (spring98):

Thanks for helping me!

OpenStudy (loser66):

yw